OpenStudy (anonymous):

Select the two values of x that are roots of the given polynomial below. x2 - 5x + 2

OpenStudy (anonymous):

(x-5/2)^2=-2+25/4=17/4 (x-5/2)=+,-square root 17/4 X= 5/2+,-square root 17/4

OpenStudy (anonymous):

i dont think those answers were right ktnguyen

OpenStudy (campbell_st):

slight error (5/2)^2 = 25/4 = 6 1/4 which is not the constant term 2...

OpenStudy (campbell_st):

use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] and if x^2 - 5x + 2 = 0 then a = 1, b = -5 and c = 2 substitute and evaluate

OpenStudy (anonymous):

thanks!!